OpenStudy (anonymous):

Evaluate 2x − y − z when x = −2, y = 3, and z = −2

OpenStudy (jackellyn):

Simply plug in the numbers they give you. So 2(-2)-(3)-(-2)=?

OpenStudy (anonymous):

-5

OpenStudy (jackellyn):

Yup!

OpenStudy (anonymous):

2(-2)-(3)-(-2)= -4-3+2= -7+2= -5